OpenStudy (anonymous):

Simplify ((k+1)!-1)+(k+1)! (k+1)

OpenStudy (anonymous):

it's part of a proof by induction question. My algebra skills always let me down =/

OpenStudy (anonymous):

I guess (k+1)! is a common factor

OpenStudy (anonymous):

confused. why is first term written \[((k+1)!-1)\] parentheses are superfluous here yes?

OpenStudy (anonymous):

can't really speculate but as it is written (k+1)! is not a common factor since it is not a factor of the first term

OpenStudy (anonymous):

-1 + (k+2)!

OpenStudy (anonymous):

yeh I can get that with wolfram alpha but how does it equate to that?

OpenStudy (anonymous):

I actually did it on paper....:-) -1 + (k+1)!(1+k+1) = -1 + (k-1)!(k+2) = -1 + (k-2)!

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

ur welcome.